OpenStudy (neycuurose1):

Solve the equation for y. (/ signs are fractions) 8x – 9y = 11 A. y= -9/8 x - 9/11 B. y = 9/8 x + 11/8 C. y= 8/9 x - 11/9 I think its between A or B...

OpenStudy (neycuurose1):

oops I meant A. y= -9/8 x - 9/11 B. y = 9/8 x + 11/8 C. y= -8x - 11 D. y=8/9 x - 11/9

OpenStudy (neycuurose1):

Can somebody please help me man.

OpenStudy (neycuurose1):

Forget it I'll just pick anything

OpenStudy (neycuurose1):